# Auxiliary Discrminator Sequence Generative Adversarial Networks (ADSeqGAN) for Few Sample Molecule Generation

### v1

@@ Abstract
In this work, we introduce Auxiliary Discriminator Sequence Generative Adversarial Networks (ADSeqGAN), a novel approach for molecular generation in small-sample datasets. Traditional generative models often struggle with limited training data, particularly in drug discovery, where molecular datasets for specific therapeutic targets, such as nucleic acids binders and central nervous system (CNS) drugs, are scarce. ADSeqGAN addresses this challenge by integrating an auxiliary random forest classifier as an additional discriminator into the GAN framework, significantly improves molecular generation quality and class specificity.
Our method incorporates pretrained generator and Wasserstein distance to enhance training stability and diversity. We evaluate ADSeqGAN on a dataset comprising nucleic acid-targeting and protein-targeting small molecules, demonstrating its superior ability to generate nucleic acid binders compared to baseline models such as SeqGAN, ORGAN, and MolGPT. Through an oversampling strategy, ADSeqGAN also significantly improves CNS drug generation, achieving a higher yield than traditional de novo models. Critical assessments, including docking simulations and molecular property analysis, confirm that ADSeqGAN-generated molecules exhibit strong binding affinities, enhanced chemical diversity, and improved synthetic feasibility.
Overall, ADSeqGAN presents a novel framework for generative molecular design in data-scarce scenarios, offering potential applications in computational drug discovery. We have demonstrated the successful applications of ADSeqGAN in generating synthetic nucleic acid-targeting and CNS drugs in this work.
@@ 1 Introduction
Non-supervised molecular generation has become a cornerstone of modern computational drug discovery, offering innovative approaches for designing novel compounds with desired properties. Over time, diverse methodologies have emerged, categorized by generative objectives and molecular representations. For instance, models have been developed for molecular property optimization, probabilistic distribution learning, and site-specific design[1]. Molecular representations vary from simplified molecular-input line-entry system (SMILES) strings and molecular graphs to molecular fingerprints and 3D point clouds[2, 3, 4]. Meanwhile, neural network architectures like Word2Vec (W2V), sequence-to-sequence (Seq2Seq) models, transformers, graph convolutional networks (GCNs), graph attention networks (GATs), message-passing neural networks (MPNNs), and 3D-point networks have powered these advancements. Generative models such as recurrent neural network (RNN), generative adversarial networks (GANs), variational autoencoders (VAEs), adversarial autoencoders (AAEs), normalizing flows, and diffusion models have further diversified the toolkit for molecular design[2, 5, 6].
@@
GANs remain a classic and versatile class of generative models, offering key advantages over VAEs and Diffusion models. By avoiding assumptions of Gaussian priors, GANs are well-suited for datasets with non-Gaussian distributions. Additionally, GANs avoid maximum likelihood estimation(MLE), which, while stabilizing optimization, can constrain generative diversity[7]. Over the years, many GAN variants have been proposed to address specific challenges in sequence generation,including Sequence GAN (SeqGAN)[8] and Objective Reinforced GAN (ORGAN)[9]. SeqGAN leverages policy gradients to optimize sequence outputs, while ORGAN incorporates task-specific rewards to guide generation through reinforcement learning (RL). On the other hand, Auxiliary Classifier GANs (ACGANs)[10] incorporate class labels into the both generator and discriminator to stabilize training, while pre-trained GAN discriminators[11], ensembles of shallow and deep classifiers, have reduced data requirements in computer vision tasks. However, such mechanisms have not yet been explored for sequence-based molecular generation tasks, leaving a gap in integrating these frameworks for SMILES-based generative models.
@@
For SMILES-based generative models, two primary objectives must be addressed during training: (1) learning the syntactic rules of SMILES notation to ensure valid molecule generation; (2) capturing the structural and functional features of molecules within the dataset. Achieving these goals often requires extensive data and carefully tuned network parameters[12, 13]. The scarcity of high-quality datasets for specific drug categories poses a significant challenge. Besides, the length of SMILES strings varies. However, many drug molecules share common features, such as small molecule binders targeting nucleic acids and proteins, which both fall under the broader category of small-molecule therapeutics. Similarly, central nervous system (CNS)-targeted drugs and other non-CNS drugs share overlapping characteristics under the drug discovery framework.
@@
Some methods have been explored to generate new molecules on small datasets based on pre-trained models and molecular scaffolds [14, 15], but the results are limited for some data sets with narrow distribution of molecular properties.
@@
In this paper, we introduce a novel approach that integrates a pretrained random forest classifier as a auxiliary descriminator into the SeqGAN framework to improve the quality of SMILES generation. To enhance adversarial training stability, we integrate MLE generator pretraining[8, 9, 16] and Wasserstein GANs (WGANs)[17, 9] into the architecture. To the best of knowledge, this GAN architechture has not been explored so far. Our method demonstrates superior performance in generating nucleic acid-targeting molecules on a nucleic acid and protein mixed dataset, achieving higher verified SMILES rates and yields for nucleic acid binders, compared to models trained exclusively on nucleic acid-targeting datasets.
@@
We also address dataset imbalances by introducing targeted data augmentation strategies. For datasets with extreme biases, such as CNS drug datasets, we employ a novel strategy of over-sampling minority molecules while training on a mixed dataset. This approach significantly increases the generation rate of CNS molecules while maintaining diversity and validity.
@@ 2 Related Work
Previous GAN-based models for SMILES sequence generation include SeqGAN and ORGAN. These foundational approaches were later extended with downstream networks such as ORGANIC[18], RANC[16] and ATNC[19], which tailored the generative process to specific application objectives.
Since the introduction of GANs[20], advancements in architectures[21, 10, 22, 23, 24], training strategies[25], and objective functions[26, 17, 27, 11] have led to significant progress. Despite their success in image-based tasks, many methods have not yet been applied to GANs for sequence generation. In this work, we integrate these advancements including ACGAN, pre-trained discriminators, and WGAN objective function into a sequence-generation GAN framework, with a particular focus on molecular sequence generation.
Molecules, due to their structured nature and extensive prior knowledge, are particularly well-suited for transfer learning. Descriptor-generation tools like RDKit[28] and OpenBabel[29] allow for the extraction of rich molecular features, which can be effectively transferred to unseen tasks, datasets, and domains. In our work, we leverage these transferable molecular property representations for unsupervised model training, enabling the generation of high-quality SMILES strings even with limited training data.
GANs can amplify data to address data scarcity issue in molecule predicting task[30], outperforming traditional methods like Synthetic Minority Oversampling Technique (SMOTE). At present, data enhancement methods have not been used to generate molecules in GANs. For highly imbalanced datasets, we employ over-sampling to enrich minority classes.
@@ 3 Methods
Figure 1: Scheme of ADSeqGAN. $Build\ Dataset$: The dataset contains 2 part, including class labels and SMILES strings. It must contains 2 or more different molecule classes. $Pretrain\ Discriminators$: Using RDKit and Open Babel to calculate molecule descriptors for every datapoints and then choose the descriptors with the strong resolution to build the classifier. To get the pretrained $D$, structural limitation is added. $Train\ ADSeqGAN$: $G$ is fed with labeled start tokens and traind by RL to generate synthetic data of different classes. $D$ with CNN is designed to ditinguish generated data from real samples, while pretrained $D$ is used to classify the samples. Reward is a combination of adversarial and auxiliary reward, and is passed back to the policy function via Monte Carlo sampling. We add length weight and repetition penalization.
@@
However, these traditional GANs can not be directly applied to sequence generation tasks, so GAN+RL based SeqGAN and ORGAN were developed. Nonetheless, in our test task, these models struggled to learn enough features with limited real data. To solve these problems, we propose Auxiliary Discrminator Sequence Generative Adversarial Networks (ADSeqGANs). The process of applying ADSeqGANs to molecular generation is shown in Figure 1: Firstly, a hybrid database consisting of a small number of samples of the desirable class and samples of auxiliary classes is constructed. Next the molecule descriptors with strong classification ability are selected by logistic regression method as parameters for pre-training to get a classifier, which is then added to GAN training as auxiliary discriminators. At the same time, labels of the real data are also entered into the generator to assist in generating samples. Last, we fine-tune the structure and parameters of the network.
Overall, this new network can benefit us in three ways: First, training shallow classifiers on pre-trained features is a common way to adapt deep networks to small data sets while reducing overfitting[31, 32, 33]; Second, the discriminator constructed based on prior physical and chemical knowledge may be more in line with human perception[34]; Third, adding standardization based on real sample or prior knowledge can reduce mode collapse[35].
@@ 3.1 Formulation
where $\text{where }D_{n}=a_{c}C_{n}(\mathcal{F})+b_{c}$
@@
where $\text{where }\sum_{n=0}^{N}\lambda_{n}=1$
@@
For discrete dataset, like SMILES strings of molecules, the sampling process is undifferentiable. One successful approach is to train $G_{\theta}$ using a RL model via policy gradient[36, 8, 9]. Considering a full length sequence $Y_{1:T}=string(t_{1}t_{2}...t_{T})$, representing a discrete data, $Y_{1:t}$ is an incomplete subsequence belonging to $Y_{1:T}$. One can maximize the long term reward in the policy gradient process to mimic the expectation in $eq(1)$:
@@ 3.2 Implementation Details
$G_{\theta}$ is a RNN with long short-term memory (LSTM) cells for sequence generation, while $D_{\phi}$ is a convolutional neural network (CNN) for text classification tasks[37]. Unlike SeqGAN and ORGAN, ADSeqGAN uses labeled start token as input, and different classes share the same RNN.
@@
To avoid problems of GAN convergence like "perfect discriminator" and improve the stability of learning, we introduce the Wasserstein-1 distance, also known as earth mover’s distance[38], to $D_{\psi}$. In essence, Wasserstein distance turns the discriminator’s classification task into a regression task, with the goal of reducing the distance between the real data distribution and the model distribution, which is more in line with the goals of RL.
For pretrained discriminator $D_{n}$, we add some structural restrictions to the classifier, in order to get more reasonable molecules.
@@
Other additional mechanisms to prevent mode collapse and over-fitting includes: (1) Using $0-1$ standardization in rollout policy when computing rewards. (2) Penalizing repeated sequences by dividing the reward of a non-unique sample by the number of copies. (3) Applying length weight to make sure the length of molecules generated correspond the real length distribution of samples.
The dataset of small molecules targeting nucleic acids and protein is built from ChEMBL345 [39, 40] and all open structures on DrugBank[41]. For the dataset of CNS and non-CNS drugs, we collect the CNS data and non-CNS data from DrugBank. We reconstruct all the SMILES strings to make them more suitable to train our models, removing the complex 3D structural information, as well as the inorganic drug, the inorganic ionic components of the drug, molecules too long, and molecules containing other than halogens and B, S, P, and canonize the sample using RDKit. All the code of tensorflow version and parameters are available online: https://github.com/ClickFF/ADSeqGAN.
@@ 4 Results
Table 1: Evaluation of metrics on several generative algorithms
Algorithm Dataset Novelty/$\text{SA}^{\Uparrow}$ $\text{QED}^{\Uparrow}$ $\text{FCD}^{\Downarrow}$ NA ratio/NA yield/Size
Validity/Unique/%
@@
Here we conduct experiments to test ADSeqGAN in two representative scenarios: a moderately biased dataset consists of 4894 nucleic acid binders (NA) and 8191 protein targeting molecules (Pro) with high bioactivity, and an extremely biased dataset consisting of 548 CNS drugs and 7728 other small molecule drugs. Our objective is to prove that our model can generate functional drug molecules with only a small number of target samples while promoting synthesizability and drug-likeness, diversity and similar molecule length distribution. To estimate the quality of generated samples, score functions designed for synthesizability (SA)[42],quantitative estimate of drug-likeness (QED)[43, 44, 45] and Frechet ChemNet Distance (FCD)[46] are used, and we use fingerprint ECFP4[47] to compute Tanimoto similarity[48] for diversity evaluation.
During training, every generator model was pretrained for 250 epochs based on MLE, and the discrminator was pre-trained for 10 steps. Unless otherwise stated, all GAN models use Wasserstein distance. Unless specified,$\lambda$ is set to 0.2. MLE-based RNN, SeqGAN, RNN-pretrained RL, chemical language model (CLM)[14], MolGPT[49], MolGen[15] are selected for contrast experiments. All GAN-related models and RL-based models are set to be pretrained for 250 epochs of MLE-based $G_{\theta}$, and then 50 epochs for downstream training. Meanwhile, in pretraining steps of GANs, $D_{\psi}$ is set to be pretrained for 10 steps. For other models, the training parameters are based on original papers. Notably, for all drug molecules in GAN and RL relative models, grammar restrictions are loose and the later generated samples are dealt with openbabel’s gen2d function to reconstruct strictly valid molecules, since lightly invalid SMILES are beneficial rather than detrimental to chemical language models[50].

### v2

@@ Abstract
In this work, we introduce Auxiliary Discriminator Sequence Generative Adversarial Networks (ADSeqGAN), a novel approach for molecular generation in small-sample datasets. Traditional generative models often struggle with limited training data, particularly in drug discovery, where molecular datasets for specific therapeutic targets, such as nucleic acids binders and central nervous system (CNS) drugs, are scarce. ADSeqGAN addresses this challenge by integrating an auxiliary random forest classifier as an additional discriminator into the GAN framework, significantly improves molecular generation quality and class specificity. Our method incorporates pretrained generator and Wasserstein distance to enhance training stability and diversity. We evaluate ADSeqGAN across three representative cases. First, on nucleic acid- and protein-targeting molecules, ADSeqGAN shows superior capability in generating nucleic acid binders compared to baseline models. Second, through oversampling, it markedly improves CNS drug generation, achieving higher yields than traditional de novo models. Third, in cannabinoid receptor type 1 (CB1) ligand design, ADSeqGAN generates novel druglike molecules, with 32.8% predicted actives surpassing hit rates of CB1-focused and general-purpose libraries when assessed by a target-specific LRIP-SF scoring function. Overall, ADSeqGAN offers a versatile framework for molecular design in data-scarce scenarios, with demonstrated applications in nucleic acid binders, CNS drugs, and CB1 ligands.
keywords:
generative adversarial networks; nucleic acids; central nervous system; computer-aided drug design; machine learning; contrast learning; small molecular drugs; drug database,
[…]
@@ 1 Introduction
Non-supervised molecular generation has become a cornerstone of modern computational drug discovery, offering innovative approaches for designing novel compounds with desired properties. Over time, diverse methodologies have emerged, categorized by generative objectives and molecular representations. For instance, models have been developed for molecular property optimization, probabilistic distribution learning, and site-specific design1. Molecular representations vary from simplified molecular-input line-entry system (SMILES) strings and molecular graphs to molecular fingerprints and 3D point clouds2, 3, 4. Meanwhile, neural network architectures like Word2Vec (W2V), sequence-to-sequence (Seq2Seq) models, transformers, graph convolutional networks (GCNs), graph attention networks (GATs), message-passing neural networks (MPNNs), and 3D-point networks have powered these advancements. Generative models such as recurrent neural network (RNN), generative adversarial networks (GANs), variational autoencoders (VAEs), adversarial autoencoders (AAEs), normalizing flows, diffusion models and large language models have further diversified the toolkit for molecular design5, 2, 6, 7, 8.
@@
GANs remain a classic and versatile class of generative models, offering key advantages over VAEs and Diffusion models. By avoiding assumptions of Gaussian priors, GANs are well-suited for datasets with non-Gaussian distributions. Additionally, GANs avoid maximum likelihood estimation(MLE), which, while stabilizing optimization, can constrain generative diversity9. Over the years, many GAN variants have been proposed to address specific challenges in sequence generation,including Sequence GAN (SeqGAN)10 and Objective Reinforced GAN (ORGAN)11. SeqGAN leverages policy gradients to optimize sequence outputs, while ORGAN incorporates task-specific rewards to guide generation through reinforcement learning (RL). On the other hand, Auxiliary Classifier GANs (ACGANs)12 incorporate class labels into the both generator and discriminator to stabilize training, while pre-trained GAN discriminators13, ensembles of shallow and deep classifiers, have reduced data requirements in computer vision tasks. However, such mechanisms have not yet been explored for sequence-based molecular generation tasks, leaving a gap in integrating these frameworks for SMILES-based generative models.
@@
For SMILES-based generative models, two primary objectives must be addressed during training: (1) learning the syntactic rules of SMILES notation to ensure valid molecule generation; (2) capturing the structural and functional features of molecules within the dataset. Achieving these goals often requires extensive data and carefully tuned network parameters14, 15. The scarcity of high-quality datasets for specific drug categories poses a significant challenge. Besides, the length of SMILES strings varies. However, many drug molecules share common features, such as small molecule binders targeting nucleic acids and proteins, which both fall under the broader category of small-molecule therapeutics. Similarly, central nervous system (CNS)-targeted drugs and other non-CNS drugs share overlapping characteristics under the drug discovery framework.
@@
Some methods have been explored to generate new molecules on small datasets based on pre-trained models and molecular scaffolds 16, 17, but the results are limited for some data sets with narrow distribution of molecular properties.
@@
In this paper, we introduce a novel approach that integrates a pretrained random forest classifier as a auxiliary descriminator into the SeqGAN framework to improve the quality of SMILES generation. To enhance adversarial training stability, we integrate MLE generator pretraining10, 11, 18 and Wasserstein GANs (WGANs)19, 11 into the architecture. To the best of knowledge, this GAN architechture has not been explored so far. Our method demonstrates superior performance in generating nucleic acid-targeting molecules on a nucleic acid and protein mixed dataset, achieving higher verified SMILES rates and yields for nucleic acid binders, compared to models trained exclusively on nucleic acid-targeting datasets.
@@
We also address dataset imbalances by introducing targeted data augmentation strategies. For datasets with extreme biases, such as CNS drug datasets, we employ a novel strategy of over-sampling minority molecules while training on a mixed dataset. This approach significantly increases the generation rate of CNS molecules while maintaining diversity and validity. To further illustrate the universality of ADSeqGAN, we conducted tests using the dataset of cannabinoid receptor type 1 receptor (CB1R) ligand. Overall, ADSeqGAN offers a versatile framework for molecular design in data-scarce scenarios.
@@ 2 Related Work
Previous GAN-based models for SMILES sequence generation include SeqGAN and ORGAN. These foundational approaches were later extended with downstream networks such as ORGANIC20, RANC18 and ATNC21, which tailored the generative process to specific application objectives.
Since the introduction of GANs22, advancements in architectures23, 12, 24, 25, 26, training strategies27, and objective functions28, 19, 29, 13, 30 have led to significant progress. Despite their success in image-based tasks, many methods have not yet been applied to GANs for sequence generation. In this work, we integrate these advancements including ACGAN, pre-trained discriminators, and WGAN objective function into a sequence-generation GAN framework, with a particular focus on molecular sequence generation.
Molecules, due to their structured nature and extensive prior knowledge, are particularly well-suited for transfer learning. Descriptor-generation tools like RDKit31 and OpenBabel32 allow for the extraction of rich molecular features, which can be effectively transferred to unseen tasks, datasets, and domains. In our work, we leverage these transferable molecular property representations for unsupervised model training, enabling the generation of high-quality SMILES strings even with limited training data.
GANs can amplify data to address data scarcity issue in molecule predicting task33, outperforming traditional methods like Synthetic Minority Oversampling Technique (SMOTE). At present, data enhancement methods have not been used to generate molecules in GANs. For highly imbalanced datasets, we employ over-sampling to enrich minority classes.
@@ 3 Methods
Figure 1: Scheme of ADSeqGAN. BuildDatasetBuild\ Dataset: The dataset contains 2 part, including class labels and SMILES strings. It must contains 2 or more different molecule classes. PretrainDiscriminatorsPretrain\ Discriminators: Using RDKit and OpenBabel to calculate molecule descriptors or fingerprints for every datapoints and then choose the descriptors with the strong resolution to build the classifier. To get the pretrained $D$, structural limitation is added. TrainADSeqGANTrain\ ADSeqGAN: $G$ is fed with labeled start tokens and traind by RL to generate synthetic data of different classes. $D$ with CNN is designed to ditinguish generated data from real samples, while pretrained $D$ is used to classify the samples. Molecules are generated via Monte Carlo sampling. Reward of each token in generated molecules is a linear combination of adversarial and auxiliary reward parameterized by $\lambda$, and is passed back to the policy gradient. We add length weight and repetition penalization to improve quality of new molecules.
@@
However, these traditional GANs can not be directly applied to sequence generation tasks, so GAN+RL based SeqGAN and ORGAN were developed. Nonetheless, in our test task, these models struggled to learn enough features with limited real data. To solve these problems, we propose Auxiliary Discrminator Sequence Generative Adversarial Networks (ADSeqGANs). The process of applying ADSeqGANs to molecular generation is shown in Figure 1: Firstly, a hybrid database consisting of a small number of samples of the desirable class and samples of auxiliary classes is constructed. Next the molecule descriptors with strong classification ability are selected by logistic regression method as parameters — or alternatively, molecular fingerprints can be directly employed — for pre-training to get a classifier, which is then added to GAN training as auxiliary discriminators. In parallel, labels of the real data are also entered into the generator to assist in generating samples. Last, we fine-tune the structure and parameters of the network.
Overall, this new network can benefit us in three ways: First, training shallow classifiers on pre-trained features is a common way to adapt deep networks to small data sets while reducing overfitting34, 35, 36; Second, the discriminator constructed based on prior physical and chemical knowledge may be more in line with human perception37; Third, adding standardization based on real sample or prior knowledge can reduce mode collapse38.
@@ 3.1 Formulation
$\text{where }D_{n}=a_{c}C_{n}(\mathcal{F})+b_{c}$
@@
$\text{where }\sum_{n=0}^{N}\lambda_{n}=1$
@@
For discrete dataset, like SMILES strings of molecules, the sampling process is undifferentiable. One successful approach is to train $G_{\theta}$ using a RL model via policy gradient39, 10, 11. Considering a full length sequence $Y_{1:T}=string(t_{1}t_{2}...t_{T})$, representing a discrete data, $Y_{1:t}$ is an incomplete subsequence belonging to $Y_{1:T}$. One can maximize the long term reward in the policy gradient process to mimic the expectation in $eq(1)$:
@@ 3.2 Implementation Details
$G_{\theta}$ is a RNN with long short-term memory (LSTM) cells for sequence generation, while $D_{\phi}$ is a convolutional neural network (CNN) for text classification tasks40. Unlike SeqGAN and ORGAN, ADSeqGAN uses labeled start token as input, and different classes share the same RNN.
@@
To avoid problems of GAN convergence like ”perfect discriminator” and improve the stability of learning, we introduce the Wasserstein-1 distance, also known as earth mover’s distance41, to $D_{\psi}$. In essence, Wasserstein distance turns the discriminator’s classification task into a regression task, with the goal of reducing the distance between the real data distribution and the model distribution, which is more in line with the goals of RL.
To guide the generator toward producing structurally reasonable and class-specific molecules, we incorporated a pretrained classifier-based discriminator $D_{n}$ as an auxiliary scoring module. First, we computed 192 molecular descriptors for all molecules using RDKit and OpenBabel, covering physicochemical, topological, and electronic properties. We then selected a subset of features showing high discriminative power, based on logistic regression and their AUC performance in 5-fold cross-validation, to train a random forest classifier using 100 trees. The resulting model was saved and deployed during generation scoring. Specifically, the probability of being correctly classified as ground truth label is given as an initial reward value.
In addition to classifier prediction, we implemented a rule-based filtering system to enforce physically motivated constraints. These include penalties for excessive nitrogen/oxygen/sulfur content, macrocyclic ring systems, and anti-aromatic rings, as well as rejection of molecules with high halogen ratios or extremely carbon-dominant structures. Conversely, positive scoring is applied to fused aromatic systems with suitable heteroatoms. The final discriminator score is computed as the product of the classifier probability and the physics-based structural score. Molecules with invalid SMILES, extreme topology, or poor drug-likeness are thereby suppressed, ensuring that generated outputs are both class-relevant and chemically meaningful.
@@
Other additional mechanisms to prevent mode collapse and over-fitting includes: (1) Using 010-1 standardization in rollout policy when computing rewards. (2) Penalizing repeated sequences by dividing the reward of a non-unique sample by the number of copies. (3) Applying length weight to make sure the length of molecules generated correspond the real length distribution of samples.
3.3 Dataset Preparation
We constructed two datasets for evaluating our model performance: (1) small molecules targeting nucleic acids or proteins, and (2) central nervous system (CNS) and non-CNS drugs.
Nucleic Acid and Protein Binder Dataset. This dataset was curated from ChEMBL345 42, 43 and all publicly available small-molecule structures on DrugBank 44. Molecules were selected based on annotations indicating their interaction with either nucleic acids or proteins.
CNS vs. Non-CNS Drug Dataset. CNS and non-CNS drugs were collected from DrugBank, including all approved and investigational small-molecule drugs with known CNS indications up to December 2024.
For both datasets, we applied the following preprocessing steps using RDKit:
  • Removal of 3D structural information and inorganic compounds.
  • Exclusion of molecules with excessive length or containing elements beyond C, H, N, O, halogens, B, S, and P.
  • Removal of salts and inorganic ionic components.
  • Canonicalization of SMILES strings for uniformity.
After preprocessing, the nucleic acid binder dataset contains 4,894 molecules, while the CNS drug dataset contains 548 CNS-targeting molecules approved and entering clinical trial before December 2024. The distributions and molecular properties of the datasets are summarized in Figure S5.
3.4 Docking
For nucleic acid binders, we performed local flexible docking of generated molecules to target nucleic acids using NLDock v1.0 45. NLDock has demonstrated superior local flexible docking accuracy over major docking tools to analyze nucleic acid and ligand interactions. First, BDS function in NLDock was used to define spherical binding‐site points from the receptor PDB. For each ligand, up to 100 conformations were generated using RDKit’s ETKDG method and converted via OpenBabel to MOL2 format. NLDock then docked these conformers to the target using the generated sphere points, retaining a maximum of 50 poses per ligand. Conformations were ranked by NLDock’s native scoring using SortConfs, and RMSD to a native ligand was calculated with obrms.
For CNS drugs, Molecular Operation Environment (MOE) 2019.0102 was used to perform molecular docking 46. MOE is a widely used commercially available docking software. Triangle matcher method based on London dG score was used to place ligands at the pocket site to generate 30 poses. For refinement, we used rigid receptor and GBVI—WSA dG score to get top 5 binding poses of each molecule.
@@ 4 Results
Table 1: Evaluation of metrics on several generative algorithms to generate nucleic acid binders
Algorithm Dataset Novelty/% $\text{SA}^{\Uparrow}$ $\text{QED}^{\Uparrow}$ $\text{FCD}^{\Downarrow}$ NA ratio/% NA yield/% Size
Validity/% Unique/%
@@
Here we conduct experiments to test ADSeqGAN in two representative scenarios: a moderately biased dataset consists of 4894 nucleic acid binders (NA) and 8191 protein targeting molecules (Pro) with high bioactivity, and an extremely biased dataset consisting of 548 CNS drugs and 7728 other small molecule drugs. Our objective is to prove that our model can generate functional drug molecules with only a small number of target samples while promoting synthesizability and drug-likeness, diversity and similar molecule length distribution. To estimate the quality of generated samples, score functions designed for synthesizability (SA)47, quantitative estimate of drug-likeness (QED)48, 49, 50 and Frechet ChemNet Distance (FCD)51 are used, and we use fingerprint ECFP452 to compute Tanimoto similarity53 for diversity evaluation.
During training, every generator model was pretrained for 250 epochs based on MLE, and the discrminator was pre-trained for 10 steps. Unless otherwise stated, all GAN models use Wasserstein distance. Unless specified, $\lambda$ is set to 0.2. MLE-based RNN, SeqGAN, RNN-pretrained RL, chemical language model (CLM)16, MolGPT54, MolGen17 are selected for contrast experiments. All GAN-related models and RL-based models are set to be pretrained for 250 epochs of MLE-based $G_{\theta}$, and then 50 epochs for downstream training. Meanwhile, in pretraining steps of GANs, $D_{\psi}$ is set to be pretrained for 10 steps. For other models, the training parameters are based on original papers. MolGen is used to generate new molecules based on the backbone of each single molecule, so 200 new molecules were generated per input. Notably, for all drug molecules in GAN and RL relative models, grammar restrictions are loose and the later generated samples are dealt with openbabel’s gen2d function to reconstruct strictly valid molecules, since lightly invalid SMILES are beneficial rather than detrimental to chemical language models55. For all generative models, the generation length was constrained to the range of 10 to 80 tokens to ensure chemical validity and relevance.